# Motivic cohomology of cyclic coverings

### v2

@@ 2.4 Vector bundles over smooth affine contractible varieties
Theorem 2.23.
If $X$ is a smooth affine variety of dimension $d=3$ over an algebraically closed field $k$, then the natural map $({c}_{1},{c}_{2},{c}_{3}):{V}_{3}(X)\xrightarrow{\cong}CH^{1}(X)\times CH^{2}(X)\times CH^{3}(X)$ is a bijection.
Proof.
See [KM, Theorem 2.1(iii)]. ∎
Theorem 2.24.
If $X$ is a smooth affine variety of dimension $d=3$ over an algebraically closed field $k$, then the natural map $({c}_{1},{c}_{2}):{V}_{2}(X)\xrightarrow{\cong}CH^{1}(X)\times CH^{2}(X)$ is a bijection.
Proof.
See [AF1, Theorem 1]. ∎
@@
If $X$ is stably $\mathbb{A}^{1}$-contractible and $d\leq 4$, then all algebraic vector bundles on $X$ are trivial.
@@
Let $X$ be a smooth affine variety of dimension $d\geq 3$ over an algebraically closed field $k$ of characteristic 00. Assume that $CH^{d}(X)\otimes_{\mathbb{Z}}\mathbb{Q}=0$. Then every vector bundle of rank $d$ has trivial direct summand of rank $1$.
@@ 3 Cyclic coverings and their motivic cohomology
In this section we discuss the motivic cohomology of cyclic coverings. As mentioned above, cyclic coverings give many examples of topologically contractible smooth affine complex varieties; we refer the reader to [Z, Section 5] for a nice survey of related results. It follows from the previous discussion that stably $\mathbb{A}^{1}$-contractible or $\mathbb{A}^{1}$-contractible smooth affine varieties have the integral motivic cohomology of the base field. In view of Conjecture 2.15 it is therefore of interest to study the motivic cohomology of cyclic coverings. We will be focusing on cyclic coverings of the following form (cf. [Z, Definition 5.1]):
Definition 3.1.
Let $k$ be a field of characteristic 00, let $X$ be an affine $k$-variety and let $f\in\mathcal{O}_{X}(X)\setminus\{0\}$ be a regular function and let $s>1$ be an integer. Then
  • we denote by $F_{0}$ the closed subscheme of $X$ defined by $f$;
  • we will denote by $F$ the closed subscheme of $Y_{s}$ defined by $u=0$.
Troughout this section, we will use the notation from the definition above. We also make the following assumptions:
General Assumptions 3.2.
Furthermore, we will assume that
  • $k$ is algebraically closed;
@@
Note that if $\mathcal{O}_{X}(X)$ is a UFD, then Eisenstein’s criterion for irreducibility implies that the third assumption holds as soons as $f$ is a prime element in $\mathcal{O}_{X}(X)$.
@@
If $U,V$ are smooth $k$-varieties and there is a Galois morphism $\varphi:U\rightarrow V$ with Galois group $G$, then $\varphi$ induces a bijection $Cor_{k}(V,W)\cong{Cor_{k}(U,W)}^{G}$ for any $W\in Sm_{k}$.
@@
But for a Galois morphism as $\varphi$ the induced morphism $U\times G\rightarrow U\times_{V}U,(u,g)\mapsto(u,u\ast g)$, where $U\times G$ denotes the disjoint union $\sqcup_{g\in G}U_{g}$ of copies of $U$, is an isomorphism. So, up to this isomorphism, one obtains an exact sequence
@@
For $g\in\mu_{s}(k)$, we will denote the automorphism $y\mapsto y\ast g$ of $Y_{s}\setminus F$ just by $g$.
@@
There is a finite correspondence $\mu:X\setminus F_{0}\rightarrow Y_{s}\setminus F$ such that $\mu\circ\varphi_{s}=\sigma:=\sum_{g\in\mu_{s}(k)}g$. Furthermore, one has $\varphi_{s}\circ\mu=s\cdot id_{X\setminus F_{0}}$.
@@
The first statement follows from the previous lemma for $\varphi=\varphi_{s}$ and $W=Y_{s}\setminus F$. For the second statement, first apply the previous lemma to $\varphi=\varphi_{s}$ and $W=X\setminus F_{0}$ in order to get the identification $Cor_{k}(X\setminus F_{0},X\setminus F_{0})\cong{Cor_{k}(Y_{s}\setminus F,X\setminus F_{0})}^{G}$ induced by $\varphi_{s}$. So it follows in particular that it suffices to show that $\varphi_{s}\circ\mu\circ\varphi_{s}=s\cdot\varphi_{s}$ for the second statement; but this follows directly from the first statement. ∎
@@
In conclusion, as $\varphi_{s}:Y_{s}\setminus F\rightarrow X\setminus F$ is a finite étale Galois cover with Galois group $\mu_{s}(k)$, it follows from th previous two paragraphs that one can define $\mu$ as the finite correspondence from $X\setminus F_{0}$ to $Y_{s}\setminus F$ represented by the transpose $\Gamma_{\varphi_{s}}^{t}$ of the graph of $\varphi_{s}$. Then this finite correspondence satisfies the properties of $\mu$ in the previous corollary. In fact, as $\varphi_{s}:Y_{s}\rightarrow X$ is at least also finite, the first paragraph above shows that the finite correspondence $\overline{\mu}$ from $X$ to $Y_{s}$ represented by the transpose $\Gamma_{\varphi_{s}}^{t}$ of the graph of $\varphi_{s}$ satisfies $\varphi_{s}\circ\overline{\mu}=s\cdot id_{X}$.
@@
Let $(p,q)\in\mathbb{Z}^{2}$. If $s\in R^{\times}$, then $\varphi_{s}$ induces isomorphisms ${H^{p,q}(Y_{s}\setminus F,R)}^{\mu_{s}(k)}\cong H^{p,q}(X\setminus F_{0},R)$. The analogous statement holds for any presheaf of $R$-modules with transfers.
@@
Follows formally from the previous corollary and representability of motivic cohomology in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ (cf. [MVW, Proposition 14.16]), which implies that motivic cohomology groups are presheaves of $R$-modules with transfers. Note that ${(\varphi_{s})}^{\ast}$ automatically maps into ${H^{p,q}(Y_{s}\setminus F,R)}^{\mu_{s}(k)}$ as $\varphi_{s}\circ g=\varphi_{s}$ and hence $g^{\ast}\circ{(\varphi_{s})}^{\ast}={(\varphi_{s})}^{\ast}$. ∎
@@
$0\rightarrow\ker(\mu^{\ast})\rightarrow{H^{p,q}(Y_{s}\setminus F,R)}\xrightarrow{\mu^{\ast}}H^{p,q}(X\setminus F_{0},R)\rightarrow 0$
and, in particular, an isomorphism ${H^{p,q}(Y_{s}\setminus F,R)}\cong H^{p,q}(X\setminus F_{0},R)\oplus\ker(\mu^{\ast})$. The analogous statements hold for any presheaf of $R$-modules with transfers.
We say that the Galois action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ is cohomologically trivial on $H^{p,q}(Y_{s}\setminus F,R)$ if ${H^{p,q}(Y_{s}\setminus F,R)}^{\mu_{s}(k)}={H^{p,q}(Y_{s}\setminus F,R)}$, i.e., any $g\in\mu_{s}(k)$ acts as identity on the group $H^{p,q}(Y_{s}\setminus F,R)$. Moreover, we say that the Galois action is motivically trivial if $g$ is the identity on $R_{tr}(Y_{s}\setminus F)$ in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ for all $g\in\mu_{s}(k)$. By representability of motivic cohomology in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ (cf. [MVW, Proposition 14.16]), it follows that motivic triviality is stronger than cohomological triviality. The following statement is a direct consequence of Corollary 3.7:
@@
Let $(p,q)\in\mathbb{Z}^{2}$ and assume $s\in R^{\times}$. Then the Galois action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ is cohomologically trivial on $H^{p,q}(Y_{s}\setminus F,R)$ if and only if $\varphi_{s}$ induces isomorphisms ${H^{p,q}(Y_{s}\setminus F,R)}\cong H^{p,q}(X\setminus F_{0},R)$. The analogous statement holds for any presheaf of $R$-modules with transfers.
@@
If $s\in R^{\times}$ and the Galois action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ is motivically trivial, then $\varphi_{s}$ induces isomorphisms $R_{tr}(Y_{s}\setminus F)\cong R_{tr}(X\setminus F_{0})$ in $\textbf{DM}_{Nis}^{eff,-}(k,R)$.
@@
Assume $s\in R^{\times}$ and let $(p,q)\in\mathbb{Z}^{2}$. Then $Y_{s}\setminus F\rightarrow Spec(k)$ induces an isomorphism on the motivic cohomology groups with coefficients in $R$ in bidegree $(p,q)$ if and only if $X\setminus F_{0}\rightarrow Spec(k)$ induces an isomorphism on the motivic cohomology groups with coefficients in $R$ in bidegree $(p,q)$ and the Galois action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ is cohomologically trivial on $H^{p,q}(Y_{s}\setminus F,R)$.
@@
Let $R$ be a commutative ring with $s\in R^{\times}$ and let $(p,q)\in\mathbb{Z}^{2}$. Then it follows in particular from the previous corollaries that the structure morphism $X\setminus F_{0}\rightarrow Spec(k)$ induces an isomorphism on motivic cohomology with coefficients in $R$ in bidegree $(p,q)$ whenever the structure morphism $Y_{s}\setminus F\rightarrow Spec(k)$ does. By using Remark 3.6 and, in particular, the finite correspondence $\overline{\mu}$, it follows that also that the structure morphism $X\rightarrow Spec(k)$ induces an isomorphism on motivic cohomology with coefficients in $R$ in bidegree $(p,q)$ whenever the structure morphism $Y_{s}\rightarrow Spec(k)$ does. In fact, it follows from Remark 3.6 that $\varphi_{s}^{\ast}:H^{p,q}(X,R)\rightarrow H^{p,q}(Y_{s},R)$ is always injective.
In view of Corollary 3.9, it is now of interest to investigate under which circumstances the Galois action of $\mu_{s}(k)$ on $H^{p,q}(Y_{s}\setminus F,R)$ is indeed trivial. We will say that an action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ extends to a morphism $h:\mathbb{A}^{1}_{k}\times_{k}(Y_{s}\setminus F)\rightarrow Y_{s}\setminus F$ (resp. $h:\mathbb{G}_{m,k}\times_{k}(Y_{s}\setminus F)\rightarrow Y_{s}\setminus F$) if there is a morphism $h:\mathbb{A}^{1}_{k}\times_{k}(Y_{s}\setminus F)\rightarrow Y_{s}\setminus F$ (resp. $h:\mathbb{G}_{m,k}\times_{k}(Y_{s}\setminus F)\rightarrow Y_{s}\setminus F$) which recovers the given action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ when evaluated at the $k$-points of $\mathbb{A}^{1}_{k}$ (resp. $\mathbb{G}_{m,k}$) corresponding to the elements of $\mu_{s}(k)$.
@@
Since any $(g,id)$ defines a section to the projection of $\mathbb{A}^{1}_{k}\times_{k}(Y_{s}\setminus F)$ onto $Y_{s}\setminus F$ and this projection is an isomorphism in $\textbf{DM}_{Nis}^{eff,-}(k,R)$, it follows that ${(g,id)}={(1,id)}$ in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ for any $g\in\mu_{s}(k)$. Hence $g$ equals the identity in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ for all $g\in\mu_{s}(k)$. ∎
@@
[…]
@@
Assume that there is a $\mathbb{G}_{m,k}$-action $\gamma$ on $X$ which makes $f$ a quasi-invariant of weight $d\in\mathbb{Z}$ with respect to $\gamma$ (cf. [Z, Definition 5.1]). One obtains a new $\mathbb{G}_{m,k}$-action $\tilde{\gamma}$ on $Y_{s}$ by setting $\lambda\ast_{\tilde{\gamma}}(x,u):=(\lambda^{s}\ast_{\gamma}x,\lambda^{d}u)$; the action restricts to an action on $Y_{s}\setminus F$. If we furthermore assume that $\langle d,s\rangle=\mathbb{Z}$, then $X$ (resp. $X\setminus F_{0}$) becomes the quotient of $Y_{s}$ (resp. $Y_{s}\setminus F$) under the restriction of $\tilde{\gamma}$ to the subscheme $\mu_{s}$ of $s$th roots of unity (over $k$), but we do not use this fact in this paper. Note that when these restrictions of $\tilde{\gamma}$ to $\mu_{s}$ are evaluated at $k$-points, i.e., actual $s$th roots of unity over $k$, then one recovers the $d$-th powers of the Galois actions of $\mu_{s}(k)$ on $X$. Since $\langle d,s\rangle=\mathbb{Z}$, taking the $d$-th power defines an automorphism of $\mu_{s}$ and we need not distinguish between the evaluation of $\tilde{\gamma}$ at $k$-points of $\mu_{s}$ and the Galois action of $\mu_{s}(k)$ when considering the groups ${H^{p,q}(Y_{s}\setminus F,R)}^{\mu_{s}(k)}$. So we can actually use the reasoning from the proof of Lemma 3.14 in order to conclude: If there is a $\mathbb{G}_{m,k}$-action $\gamma$ on $X$, which makes $f$ a quasi-invariant of weight $d\in\mathbb{Z}$ with respect to $\gamma$ and $\langle d,s\rangle=\mathbb{Z}$, then the Galois action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ is cohomologically trivial on $H^{p,q}(Y_{s}\setminus F,R)$ as soon as the projection $\pi:\mathbb{G}_{m,k}\times_{k}(Y_{s}\setminus F)\rightarrow(Y_{s}\setminus F)$ induces an isomorphism $\pi^{\ast}:H^{p,q}(Y_{s}\setminus F,R)\xrightarrow{\cong}H^{p,q}(\mathbb{G}_{m,k}\times_{k}Y_{s}\setminus F,R)$.
Remark 3.17.
The $\mathbb{G}_{m,k}$-action $\tilde{\gamma}$ in the beginning of Remark 3.16 above also shows: If there is a $\mathbb{G}_{m,k}$-action $\gamma$ on $X$ which makes $f$ a quasi-invariant of weight $d\in\mathbb{Z}$ with respect to $\gamma$, then there is also a $\mathbb{G}_{m,k}$-action $\tilde{\gamma}$ on $X\times_{k}\mathbb{A}^{1}_{k}$ which makes $f-u^{s}$ (where $u$ is the coordinate of $\mathbb{A}^{1}_{k}$) a quasi-invariant of weight $ds\in\mathbb{Z}$ with respect to $\tilde{\gamma}$.

### v3

@@ 2.4 Vector bundles over smooth affine contractible varieties
Theorem 2.23 ([KM, Theorem 2.1(iii)]).
If $X$ is a smooth affine variety of dimension $d=3$ over an algebraically closed field $k$ of characteristic $\neq 2$, then the natural map $({c}_{1},{c}_{2},{c}_{3}):{V}_{3}(X)\xrightarrow{\cong}CH^{1}(X)\times CH^{2}(X)\times CH^{3}(X)$ is a bijection.
Theorem 2.24 ([AF1, Theorem 1]).
If $X$ is a smooth affine variety of dimension $d=3$ over an algebraically closed field $k$ of characteristic $\neq 2$, then the natural map $({c}_{1},{c}_{2}):{V}_{2}(X)\xrightarrow{\cong}CH^{1}(X)\times CH^{2}(X)$ is a bijection.
@@
If $X$ is a stably $\mathbb{A}^{1}$-contractible smooth affine variety of dimension $d\leq 4$ over an algebraically closed field with $6\in k^{\times}$, then all algebraic vector bundles on $X$ are trivial.
@@
Let $X$ be a smooth affine variety of dimension $d\geq 3$ over an algebraically closed field $k$ of characteristic 0. Assume that $CH^{d}(X)\otimes_{\mathbb{Z}}\mathbb{Q}=0$. Then every vector bundle of rank $d$ has trivial direct summand of rank $1$.
@@ 3 Cyclic coverings and their motivic cohomology
In this section we discuss the motivic cohomology of cyclic coverings. As mentioned above, cyclic coverings give many examples of topologically contractible smooth affine complex varieties; we refer the reader to [Z, Section 5] for a nice survey of related results. It follows from the previous discussion that stably $\mathbb{A}^{1}$-contractible or $\mathbb{A}^{1}$-contractible smooth affine varieties have the integral motivic cohomology of the base field. In view of Conjecture 2.15 it is therefore of interest to study the motivic cohomology of cyclic coverings. We will be focusing on cyclic coverings $\varphi_{s}:Y_{s}\rightarrow X$ as defined in Definition 1.1. Troughout this section, we will use the notation from Definition 1.1 and work under the General Assumptions 1.2.
@@
Note that if $\mathcal{O}_{X}(X)$ is a UFD, then Eisenstein’s criterion for irreducibility implies that the second assumption in the General Assumptions 1.2 holds as soons as $f$ is a prime element in $\mathcal{O}_{X}(X)$.
@@
We will now prove several statements about finite étale Galois covers which, in particular, apply to the morphism $\varphi_{s}:Y_{s}\setminus F\rightarrow X\setminus F_{0}$ above. For this purpose, we fix a finite étale Galois cover $\varphi:U\rightarrow V$ between smooth $k$-varieties with Galois group $G$ and of degree $s=|G|$. We will call the corresponding action of the group $G$ on $U$ the Galois action and denote it usually by $\ast$.
@@
Let $\varphi:U\rightarrow V$ be a finite étale Galois cover between smooth $k$-varieties with Galois group $G$. Then $\varphi$ induces a bijection $Cor_{k}(V,W)\cong{Cor_{k}(U,W)}^{G}$ for any $W\in Sm_{k}$.
@@
But for a finite étale Galois cover as $\varphi$ the induced morphism $U\times G\rightarrow U\times_{V}U,(u,g)\mapsto(u,u\ast g)$, where $U\times G$ denotes the disjoint union $\sqcup_{g\in G}U_{g}$ of copies of $U$, is an isomorphism. So, up to this isomorphism, one obtains an exact sequence
@@
For $g\in G$, we will denote the automorphism $y\mapsto y\ast g$ of $U$ just by $g$.
@@
There is a finite correspondence $\mu:V\rightarrow U$ such that $\mu\circ\varphi=\sigma:=\sum_{g\in G}g$. Furthermore, one has $\varphi\circ\mu=s\cdot id_{X\setminus F_{0}}$.
@@
The first statement follows from the previous lemma for $\varphi$ and $W=U$. For the second statement, first apply the previous lemma to $\varphi$ and $W=V$ in order to get the identification $Cor_{k}(V,V)\cong{Cor_{k}(U,V)}^{G}$ induced by $\varphi$. So it follows in particular that it suffices to show that $\varphi\circ\mu\circ\varphi=s\cdot\varphi$ for the second statement; but this follows directly from the first statement. ∎
@@
In conclusion, as $\varphi_{s}:Y_{s}\setminus F\rightarrow X\setminus F$ is a finite étale Galois cover with Galois group $\mu_{s}(k)$, it follows from the previous two paragraphs that one can define $\mu$ as the finite correspondence from $X\setminus F_{0}$ to $Y_{s}\setminus F$ represented by the transpose $\Gamma_{\varphi_{s}}^{t}$ of the graph of $\varphi_{s}$. Then this finite correspondence satisfies the properties of $\mu$ in the previous corollary. In fact, as $\varphi_{s}:Y_{s}\rightarrow X$ is at least also finite, the first paragraph above shows that the finite correspondence $\overline{\mu}$ from $X$ to $Y_{s}$ represented by the transpose $\Gamma_{\varphi_{s}}^{t}$ of the graph of $\varphi_{s}$ satisfies $\varphi_{s}\circ\overline{\mu}=s\cdot id_{X}$.
@@
Let $(p,q)\in\mathbb{Z}^{2}$. If $s\in R^{\times}$, then $\varphi$ induces isomorphisms ${H^{p,q}(U,R)}^{G}\cong H^{p,q}(V,R)$. The analogous statement holds for any presheaf of $R$-modules with transfers.
@@
The statement follows formally from the Corollary 3.3 and representability of motivic cohomology in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ (cf. [MVW, Proposition 14.16]), which implies that motivic cohomology groups are presheaves of $R$-modules with transfers: Note that $\varphi$$^{\ast}$ automatically maps into the group ${H^{p,q}(U,R)}^{G}$ as $\varphi\circ g=\varphi$ and hence $g^{\ast}\circ{\varphi}^{\ast}={\varphi}^{\ast}$. If we let $\mu:V\rightarrow U$ be as in Corollary 3.3, then the composite
$H^{p,q}(V,R)\xrightarrow{\varphi^{\ast}}{H^{p,q}(U,R)}^{G}\hookrightarrow H^{p,q}(U,R)\xrightarrow{\mu^{\ast}}H^{p,q}(V,R)$
is multiplication by $s$ and hence, as $s\in R^{\times}$, an isomorphism; in particular, $\varphi^{\ast}:H^{p,q}(V,R)\rightarrow{H^{p,q}(U,R)}^{G}$ is injective. Similarly, the composite
${H^{p,q}(U,R)}^{G}\hookrightarrow H^{p,q}(U,R)\xrightarrow{\mu^{\ast}}H^{p,q}(V,R)\xrightarrow{\varphi^{\ast}}{H^{p,q}(U,R)}^{G}$
is $\sum_{g\in G}g^{\ast}$, which is multiplication by $s$ on ${H^{p,q}(U,R)}^{G}$ and hence an isomorphism as well because $s\in R^{\times}$; in particular $\varphi^{\ast}$ is also surjective and hence an isomorphism. The reasoning for any presheaf of $R$-modules with transfers is completely analogous. ∎
@@
$0\rightarrow\ker(\mu^{\ast})\rightarrow{H^{p,q}(U,R)}\xrightarrow{\mu^{\ast}}H^{p,q}(V,R)\rightarrow 0$
and, in particular, an isomorphism ${H^{p,q}(U,R)}\cong H^{p,q}(V,R)\oplus\ker(\mu^{\ast})$. The analogous statements hold for any presheaf of $R$-modules with transfers.
We say that the Galois action of $G$ on $U$ is cohomologically trivial on $H^{p,q}(U,R)$ if ${H^{p,q}(U,R)}^{G}={H^{p,q}(U,R)}$, i.e., any $g\in G$ acts as identity on the group $H^{p,q}(U,R)$. Moreover, we say that the Galois action is motivically trivial if $g$ is the identity on $R_{tr}(U)$ in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ for all $g\in G$. By representability of motivic cohomology in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ (cf. [MVW, Proposition 14.16]), it follows that motivic triviality is stronger than cohomological triviality. The following statement is a direct consequence of Corollary 3.5:
@@
Let $(p,q)\in\mathbb{Z}^{2}$ and assume $s\in R^{\times}$. Then the Galois action of $G$ on $U$ is cohomologically trivial on $H^{p,q}(U,R)$ if and only if $\varphi$ induces isomorphisms ${H^{p,q}(U,R)}\cong H^{p,q}(V,R)$. The analogous statement holds for any presheaf of $R$-modules with transfers.
@@
If $s\in R^{\times}$ and the Galois action of $G$ on $U$ is motivically trivial, then $\varphi$ induces isomorphisms $R_{tr}(U)\cong R_{tr}(V)$ in $\textbf{DM}_{Nis}^{eff,-}(k,R)$.
@@
Assume $s\in R^{\times}$ and let $(p,q)\in\mathbb{Z}^{2}$. Then $U\rightarrow Spec(k)$ induces an isomorphism on the motivic cohomology groups with coefficients in $R$ in bidegree $(p,q)$ if and only if $V\rightarrow Spec(k)$ induces an isomorphism on the motivic cohomology groups with coefficients in $R$ in bidegree $(p,q)$ and the Galois action of $G$ on $U$ is cohomologically trivial on $H^{p,q}(U,R)$.
@@
Let $R$ be a commutative ring with $s\in R^{\times}$ and let $(p,q)\in\mathbb{Z}^{2}$. Then it follows in particular from the previous corollaries that the structure morphism $X\setminus F_{0}\rightarrow Spec(k)$ induces an isomorphism on motivic cohomology with coefficients in $R$ in bidegree $(p,q)$ whenever the structure morphism $Y_{s}\setminus F\rightarrow Spec(k)$ does. By using Remark 3.4 and, in particular, the finite correspondence $\overline{\mu}$, it follows also that the structure morphism $X\rightarrow Spec(k)$ induces an isomorphism on motivic cohomology with coefficients in $R$ in bidegree $(p,q)$ whenever the structure morphism $Y_{s}\rightarrow Spec(k)$ does. In fact, it follows from Remark 3.4 that $\varphi_{s}^{\ast}:H^{p,q}(X,R)\rightarrow H^{p,q}(Y_{s},R)$ is always injective.
We now focus on the morphism $\varphi_{s}:Y_{s}\setminus F\rightarrow X\setminus F_{0}$. In view of Corollary 3.7, it is now of interest to investigate under which circumstances the Galois action of $\mu_{s}(k)$ on $H^{p,q}(Y_{s}\setminus F,R)$ is indeed trivial. We will say that an action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ extends to a morphism $h:\mathbb{A}^{1}_{k}\times_{k}(Y_{s}\setminus F)\rightarrow Y_{s}\setminus F$ (resp. $h:\mathbb{G}_{m,k}\times_{k}(Y_{s}\setminus F)\rightarrow Y_{s}\setminus F$) if there is a morphism $h:\mathbb{A}^{1}_{k}\times_{k}(Y_{s}\setminus F)\rightarrow Y_{s}\setminus F$ (resp. $h:\mathbb{G}_{m,k}\times_{k}(Y_{s}\setminus F)\rightarrow Y_{s}\setminus F$) which recovers the given action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ when evaluated at the $k$-points of $\mathbb{A}^{1}_{k}$ (resp. $\mathbb{G}_{m,k}$) corresponding to the elements of $\mu_{s}(k)$.
@@
Since any $(g,id)$ defines a section to the projection of $\mathbb{A}^{1}_{k}\times_{k}(Y_{s}\setminus F)$ onto $Y_{s}\setminus F$ and this projection is an isomorphism in $\textbf{DM}_{Nis}^{eff,-}(k,R)$, it follows that (g,id)=(1,id){(g,id)}={(1,id)} in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ for any $g\in\mu_{s}(k)$. Hence $g$ equals the identity in $\textbf{DM}_{Nis}^{eff,-}(k,R)$ for all $g\in\mu_{s}(k)$. ∎
@@
$Y_{s}\setminus F\xrightarrow{(id,1)}(Y_{s}\setminus F)\times_{k}\mathbb{G}_{m,k}\xrightarrow{\pi}Y_{s}\setminus F$
@@
Lemma 3.14.
Let $(p,q)\in\mathbb{Z}^{2}$. Assume that there is a $\mathbb{G}_{m,k}$-action $\gamma$ on $X$ which makes $f$ a quasi-invariant of weight $d\in\mathbb{Z}$ with respect to $\gamma$ and $\langle d,s\rangle=\mathbb{Z}$. Then the Galois action of $\mu_{s}(k)$ on $Y_{s}\setminus F$ is cohomologically trivial on $H^{p,q}(Y_{s}\setminus F,R)$ as soon as the projection $\pi:\mathbb{G}_{m,k}\times_{k}(Y_{s}\setminus F)\rightarrow(Y_{s}\setminus F)$ induces an isomorphism $\pi^{\ast}:H^{p,q}(Y_{s}\setminus F,R)\xrightarrow{\cong}H^{p,q}(\mathbb{G}_{m,k}\times_{k}Y_{s}\setminus F,R)$.
Proof.
By assumption, there is a $\mathbb{G}_{m,k}$-action
$\gamma:\mathbb{G}_{m,k}\times_{k}X\rightarrow X,(\lambda,x)\mapsto\lambda\ast_{\gamma}x$
on $X$ which makes $f$ a quasi-invariant of weight $d\in\mathbb{Z}$ with respect to $\gamma$. As explained after [Z, Definition 5.1], one obtains a new $\mathbb{G}_{m,k}$-action $\tilde{\gamma}$ on $Y_{s}$ by setting
$\lambda\ast_{\tilde{\gamma}}(x,u):=(\lambda^{s}\ast_{\gamma}x,\lambda^{d}u)$.
More precisely, let
${()}^{l}:\mathbb{G}_{m,k}\rightarrow\mathbb{G}_{m,k}$
denote $l$th power morphism on $\mathbb{G}_{m,k}$ for $l\in\mathbb{Z}$. Then we first obtain a new $\mathbb{G}_{m,k}$-action $\gamma_{1}$ on $X$ via the composite
$\gamma_{1}:\mathbb{G}_{m,k}\times_{k}X\xrightarrow{{()}^{s}\times_{k}id_{X}}\mathbb{G}_{m,k}\times_{k}X\xrightarrow{\gamma}X$.
Similarly, if we consider the multiplication action of $\mathbb{G}_{m,k}$ on $\mathbb{A}^{1}_{k}$ (with coordinate $u$)
$m:\mathbb{G}_{m,k}\times_{k}\mathbb{A}^{1}_{k}\rightarrow\mathbb{A}^{1}_{k},(\lambda,u)\mapsto\lambda\cdot u$,
we obtain a new $\mathbb{G}_{m,k}$-action $\gamma_{2}$ on $\mathbb{A}^{1}_{k}$ via the composite
$\gamma_{2}:\mathbb{G}_{m,k}\times_{k}\mathbb{A}^{1}_{k}\xrightarrow{{()}^{d}\times_{k}id_{\mathbb{A}^{1}_{k}}}\mathbb{G}_{m,k}\times_{k}\mathbb{A}^{1}_{k}\xrightarrow{m}\mathbb{A}^{1}_{k}$.
If we let $\pi_{1}:\mathbb{G}_{m,k}\times_{k}X\times_{k}\mathbb{A}^{1}_{k}\rightarrow\mathbb{G}_{m,k}\times_{k}X$ and $\pi_{2}:\mathbb{G}_{m,k}\times_{k}X\times_{k}\mathbb{A}^{1}_{k}\rightarrow\mathbb{G}_{m,k}\times_{k}\mathbb{A}^{1}_{k}$ denote the projections, we then obtain an action $\tilde{\gamma}$ on $X\times_{k}\mathbb{A}^{1}_{k}$ which is simply defined as the product
$\mathbb{G}_{m,k}\times_{k}X\times_{k}\mathbb{A}^{1}_{k}\xrightarrow{(\gamma_{1}\circ\pi_{1},\gamma_{2}\circ\pi_{2})}X\times_{k}\mathbb{A}^{1}_{k}$.
By construction, the action $\tilde{\gamma}$ makes the function $f-u^{s}\in\mathcal{O}_{X\times_{k}\mathbb{A}^{1}_{k}}(X\times_{k}\mathbb{A}^{1}_{k})=\mathcal{O}_{X}(X)[u]$ a quasi-invariant of weight $ds\in\mathbb{Z}$; in particular, $\tilde{\gamma}$ descends to a $\mathbb{G}_{m,k}$-action on $Y_{s}$. By abuse of notation, we also call this action $\tilde{\gamma}$ and this is precisely the action on $Y_{s}$ indicated in the first paragraph of this proof. The action $\tilde{\gamma}$ on $Y_{s}$ clearly restricts to an action on $Y_{s}\setminus F$ (which, by abuse of notation, we also call $\tilde{\gamma}$).
Now we use the assumption that $\langle d,s\rangle=\mathbb{Z}$. Note that when the restriction of $\tilde{\gamma}$ to $\mu_{s}$ is evaluated at $k$-points, i.e., actual $s$th roots of unity over $k$, then one recovers the $d$-th powers of the Galois actions of $\mu_{s}(k)$ on $X$. Since $\langle d,s\rangle=\mathbb{Z}$, taking the $d$-th power defines an automorphism of $\mu_{s}$ and we need not distinguish between the evaluation of $\tilde{\gamma}$ at $k$-points of $\mu_{s}$ and the Galois action of $\mu_{s}(k)$ when considering the groups ${H^{p,q}(Y_{s}\setminus F,R)}^{\mu_{s}(k)}$. Now the reasoning from the proof of Lemma 3.12 (or rather Lemma 3.11) finishes the proof. ∎
@@
The $\mathbb{G}_{m,k}$-action $\tilde{\gamma}$ in the proof of Lemma 3.14 above also shows: If there is a $\mathbb{G}_{m,k}$-action $\gamma$ on $X$ which makes $f$ a quasi-invariant of weight $d\in\mathbb{Z}$ with respect to $\gamma$, then there is also a $\mathbb{G}_{m,k}$-action $\tilde{\gamma}$ on $X\times_{k}\mathbb{A}^{1}_{k}$ which makes $f-u^{s}$ (where $u$ is the coordinate of $\mathbb{A}^{1}_{k}$) a quasi-invariant of weight $ds\in\mathbb{Z}$ with respect to $\tilde{\gamma}$.